OpenStudy (anonymous):

solve by factoring the trinomial as a binomial square and then use the square root method. Round to nearest tenth if necessary. x^2 + 16x + 64 = 81

OpenStudy (anonymous):

Please help!!!! :/

OpenStudy (anonymous):

\[x^2+16x + 64 - 81 = 0\] \[x^2 + 16x - 17 = 0 \implies x^2 + 17x - x - 17 = 0\] \[x(x+17) -1(x+17) = 0 \implies (x+17)(x-1) = 0\] \[x = 1 \quad Or \quad x = -17\]

OpenStudy (anonymous):

k thanks!!

OpenStudy (anonymous):

Welcome dear...